OpenStudy (jamaljunior):

Solve for x. 1+|2+x|=9

OpenStudy (jamaljunior):

@zepdrix @mathmale plz help me

OpenStudy (mathmale):

You could greatly simplify this problem if you'd please subtract 1 from both sides. Try to explain in your own words why you'd do that. What's the next step? The resulting absolute value equation has 2 solutions. How would you go about finding them?

OpenStudy (jamaljunior):

to separate the variable?

OpenStudy (mathmale):

To simplify the equation, the whole equation. You end up with |2+x|=8. What next?

OpenStudy (jamaljunior):

divide 2/2 and 8/2?

OpenStudy (mathmale):

Why divide? There's no multiplication in |x+2|=8.

OpenStudy (jamaljunior):

because thats what i usually do when solving for x

OpenStudy (mathmale):

Unfortunately that's not good enough a reason. What is the absolute value of -3? What is the absolute value of 7?

OpenStudy (jamaljunior):

3 and 7

OpenStudy (mathmale):

Yes. The absolute value operator changes - numbers to + ones. Thus, |x+2|=8 can be broken into 2 separate equations: +(x+2)=8 and -(x-2)=8 Can you solve the first equation for x?

OpenStudy (jamaljunior):

4

OpenStudy (mathmale):

show your work, please. Unfortunately, 4 is not the solution. Simplify +(x+2)=8 now. Show your work and your result.

OpenStudy (mathmale):

You may be confusing addition with multiplication.

OpenStudy (jamaljunior):

2/2 8/2

OpenStudy (mathmale):

No division either, Jamal. Why divide here? +(x+2)=8 You can simply drop the first + sign and drop the parentheses ( ). Please do that now.

OpenStudy (jamaljunior):

10

OpenStudy (mathmale):

Jamal, did you follow my instructions? You have +(x+2)=8. I asked you to OMIT (get rid of) the first + sign. Do that now, please.

OpenStudy (jamaljunior):

x+2=8

OpenStudy (mathmale):

much better. Thank you. Now solve that for x. Subtract 2 from both sides. Show your work.

OpenStudy (jamaljunior):

2-2 = 0 8-2 = 6

OpenStudy (jamaljunior):

6x

OpenStudy (jamaljunior):

x = 6

OpenStudy (mathmale):

x=6 is correct. But your "2-2 = 0 8-2 = 6" while "right," is hard to follow / read. I recommend you do it like this: x + 2 = 8 -2 = -2 ------------ You end up with x = 6.

OpenStudy (jamaljunior):

ok thanku

OpenStudy (mathmale):

Now look at the other solution: -(x+2)=8 Think: What do you have to do first? Goal: solve this equation for x.

OpenStudy (mathmale):

You want to isolate (solve for) x.

OpenStudy (jamaljunior):

drop the - and () make it x + 2 = 8 x + 2 = 8 -2 = -2 x = 6

OpenStudy (mathmale):

Can't drop the -. -(x+2)=8 Try multiplying both sides of this equation by (-1). Show your results.

OpenStudy (jamaljunior):

x + -2 = -8

OpenStudy (mathmale):

If you have -(x+2)=8 and mult. both sides by -1, you get x+2=-8. Is this clear? If not, ask questions.

OpenStudy (jamaljunior):

why is it not -2 if you multiply both sides

OpenStudy (mathmale):

Because multiplying -(x+2) by -1 results in +(x+2).

OpenStudy (mathmale):

You still have a neg. sign in the equation, but now it's in front of 8.

OpenStudy (mathmale):